OpenStudy (pokemon23):

The width of a rectangle is 8 feet less than its length. If the perimeter is 20 feet find the measures of its length and width. The width of a rectangle is 8 feet less than its length. If the perimeter is 20 feet find the measures of its length and width. @Mathematics

OpenStudy (pokemon23):

BATH DUDE ALRIGHT

OpenStudy (anonymous):

YAYZ

OpenStudy (pokemon23):

4 people 4 medals

OpenStudy (anonymous):

w=16 - 20 cuz ther's 2 sides that are the same

OpenStudy (anonymous):

srry 20-16

OpenStudy (anonymous):

length is x width is x-8 so 2(x)+2(x-8)=20 2x+2x-16=20 4x=36 x=9

OpenStudy (anonymous):

i am right!!!

OpenStudy (pokemon23):

why is everyone getting different answers

OpenStudy (anonymous):

cuz im the only one who's right

OpenStudy (pokemon23):

so the answer is?

OpenStudy (anonymous):

nick is right

jhonyy9 (jhonyy9):

20=2l+2(l-8) 20=2l+2l-16 20+16=4l 36=4l l=36/4 l=9 w=(9-8) w=1

OpenStudy (pokemon23):

is w=1 or 9?

jhonyy9 (jhonyy9):

P=9+9+1+1=20

jhonyy9 (jhonyy9):

9 is the lenght

jhonyy9 (jhonyy9):

the width is 1